The following has been provided:

Sales $1,974,000

Direct Materials Inventory. 1/1 54,000

Direct Materials Inventory, 12/31 ?

Direct Materials Purchased 206,000

Direct Materials used in Production 212,000

Direct Labor 434,000

Manufacturing Overhead 650,000

Selling and Administrative Expenses 300,000

Costs of goods manufactured 1,322,000

Cost of goods available for sale 1,420,000

Work in Process, 1/1 156,000

Work in Process, 12/31 ?

Finished Goods 1/1 ?

Finished Goods, 12/31 116,000

Income before taxes 370,000

Required:

Determine the following amounts.

a. Direct Materials Inventory, 12/31

b. Work in Process, 12/31

c. Finished Goods, 1/1
Answer #1

a.

Direct Materials Inventory. 1/1= $54,000

Direct Materials Inventory, 12/31 = ?

Direct Materials Purchased = $206,000

Direct Materials used in Production = $212,000

Direct Materials used in Production = Direct Materials Inventory. 1/1+ Direct Materials Purchased- Direct Materials Inventory, 12/31

212,000 = 54,000+206,000-Direct Materials Inventory, 12/31

Direct Materials Inventory, 12/31 = $48,000

b.

Direct Materials used in Production = $212,000

Direct Labor= $434,000

Manufacturing Overhead = $650,000

Total manufacturing cost = Direct Materials used in Production+ Direct Labor+ Manufacturing Overhead

= 212,000+434,000+650,000

= $1,296,000

Costs of goods manufactured = $1,322,000

Work in Process, 1/1 = $156,000

Cost of goods manufactured = Total manufacturing cost+ Work in Process, 1/1- Work in Process, 12/31

1,322,000 = 1,296,000+156,000-Work in Process, 12/31

Work in Process, 12/31 = $130,000

c.

Cost of goods available for sale = $1,420,000

Cost of goods available for sale = Cost of goods manufactured +  Finished Goods, 1/1

1,420,000= 1,322,000+ Finished Goods, 1/1

Finished Goods, 1/1 = $98,000

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!